# Representing local protein environments with atomistic foundation models

### v2

@@ B Isolated $\beta$-hridge
Appendix C Extended Background
@@
Chemical shifts. The resonance frequency of a nuclear spin, $\nu$, depends directly on the splitting of its energy levels in a magnetic field. This is given as $\nu\propto\gamma B$, where $B$ is the strength of the magnetic field that the nucleus experiences and $\gamma$ is the gyromagnetic ratio of the nucleus, a characteristic of an atom’s nucleus (e.g., 1H, 13C). In Nuclear Magnetic Resonance (NMR) experiments, $B$ is primarily determined by the externally applied magnetic field $B_{0}$, but at the level of an observed atom it is slightly perturbed by local magnetic fields $\Delta B$ induced by the surrounding electron cloud. These perturbations, which are typically several orders of magnitude weaker than $B_{0}$, cause small shifts in the resonance frequency—known as chemical shifts. The chemical shift reflects subtle variations in the local electronic environment and is therefore highly informative about molecular structure and composition. The chemical change of an atom is often reported in relative terms as $\delta=10^{6}\cdot\frac{\nu-\nu_{0}}{\nu_{0}}$, where $\nu$ is the observed resonance frequency and $\nu_{0}$ is the resonance frequency of the same type of nucleus in a reference compound – commonly the 1H signal of 2,2-dimethyl-2-silapentane-5-sulfonate (DSS) is used for the 1H chemical-shift scale. Although dimensionless, chemical shifts $\delta$ are conventionally expressed in parts per million (ppm), hence the $10^{6}$ scaling factor in the definition.
@@
The electronic environment that gives rise to chemical shifts is influenced by factors such as the nature of chemical bonds, the identities of neighboring atoms, and their three-dimensional spatial arrangement. Additionally, dihedral angles [^2] The torsion angle about the bond between atoms B and C in a four-atom segment A–B–C–D play a significant role in modulating electron density around nuclei and thereby affect chemical shifts. Other factors that influence the chemical shift include hydrogen bonding interactions, ligand binding, or proximity to solvent molecules. Therefore, chemical shifts provide site-specific information about the local environment at the atomic-level. The chemical shift reports the time-averaged environment over all motions on time scales shorter than milliseconds. Chemical shifts are particularly valuable for inferring atomic connectivity, identifying functional groups, and detecting conformational changes, due to their sensitivity to fine variations in electronic structure [12, 23]. Accurate chemical shift prediction plays a critical role in applications like automated NMR resonance assignment, molecular structure determination and validation, and the analysis of complex chemical systems [68, 7]. Despite this, reliable chemical shift prediction remains a challenge due to the sensitivity of chemical shifts to nuanced changes in molecular environment [37, 9].
@@
ESM-based descriptors [43]. Evolutionary Scale Modeling (ESM) is a family of transformer-based protein language models trained on a large-scale dataset of protein amino acid sequences using masked language modeling, similar to BERT [14] for natural language. Unlike traditional methods that rely on multiple sequence alignments (MSAs) or co-evolutionary profiles, ESM models learn context-aware representations directly from amino acid sequences, enabling them to capture biochemical, structural, and evolutionary patterns within a protein sequence without explicit alignment.
@@
MACE and Egret descriptors [3, 67]. Message passing neural networks [21, 59] typically exchange messages between pairs of nodes in a graph. In contrast, the MACE family of models (MACE, Egret) generalizes this framework by learning equivariant messages involving a system of $n$ (order parameter) nodes (or atoms) at a time. This enables richer geometric and chemical modeling by operating directly over higher-order interactions. Formally, for a given node at layer $t$ of the graph, MACE first learns the following $S_{n}$-invariant and $O(3)$-equivariant features between pairs of atoms:
@@
AIMNet2-based descriptors [71]. AIMNet constructs Atomic Feature Vectors (AFVs) by iteratively embedding the local atomic environments and updating atomic representations using geometric information. Given the atomic coordinates $\mathbf{R}$ and atomic number $Z$, an initial atomic feature vector $\mathbf{A}_{i}^{(0)}$ is derived for atom $i$. The initial embeddings are parametrized as learnable vectors that depend on atomic number $Z_{i}$ of each atom.
@@
This process is repeated over multiple iterations, enabling information to propagate beyond the local neighborhood and capture non-local effects such as charge redistribution and polarization. After several refinement steps, AIMNet produces AFVs of fixed dimensionality ($256$), which are used to predict various molecular and atomic properties. Training is performed in a multitask setting, where the network jointly learns to predict a variety of molecular and atomic properties—including energies, forces, partial charges, and dipole moments—using shared representations to improve generalization. The model is trained on large datasets of quantum mechanical calculations, with regularization techniques to ensure stability and prevent overfitting.
OrbNet-based descriptors [53]. In similar spirit to semi-empirical molecular force-fields, OrbNet constructs feature vectors by leveraging symmetry-adapted atomic orbitals (AO), which capture local electronic structure information obtained from quantum calculations while respecting the symmetries inherent in molecular systems. Using mean-field quantum mechanics, OrbNet derives overlap matrices $\mathbf{S}$, Fock matrix $\mathbf{F}$, density matrix $\mathbf{D}$, and the Hamiltonian $\mathbf{H}$. Each matrix element $O^{\nu\mu}$ corresponds to interactions between atomic orbitals $\nu$ and $\mu$, and blocks $\mathbf{O}_{ij}$ represent interactions between atoms $i$ and $j$.
@@
LoCoHD [18] is a method for quantifying chemical and structural differences between protein environments. Unlike alignment-based or purely geometric measures, LoCoHD characterizes each local environment as a distribution of chemical “primitive types” such as atom types or residue centroids within a specified radius. The similarity between two environments is then measured using the Hellinger distance between their respective distributions. In this paper, to test the effectiveness of LoCoHD descriptors in representing local environments, we construct LocoHD embeddings using its representation of a local environment while computing the similarity metric. Each embedding is computed by aggregating statistics of the primitive types, weighted by their distance from the central residue. This results in a purely structural and chemical descriptor of the local environment.
Appendix D Data curation
The dataset used within the scope of this paper was sourced from RefDB [69] which is a subset of the Biomolecular Magnetic Resonance Databank (BMRB) [26]. RefDB is a curated list of BMRB entries with calibrated chemical shifts. We select monomers from the RefDB and perform sequence redundancy filtering with mmseqs via,
@@ Appendix D Data curation
which results in sequence clusters. We then split the clusters to define train and test sets. Our final dataset consists of $1048$ BMRB entries in RefDB ($225$ test + $823$ train), from where we collected the amino acid sequences, and the experimental chemical shifts. The structures are obtained by running AlphaFold2 [29] via OpenFold [2]. Using AlphaFold offers two advantages: (i) it avoids structural inconsistencies that may be inherent in crystal structures, and (ii) it allows predicting shifts for structures that do not have a crystal structure deposited to the PDB. The local protein environments are obtained from each protein structure are extracted as described in Algorithm 1. We obtain the random-coil chemical shifts used in our shift predictor from UCBShift2 [52].
Appendix E Environment construction and AFM embedding
This section elaborates on the amino-acid centric environment construction algorithm discussed in Section 3.
Algorithm 1 Amino-Acid Centric Environment Extraction
1:Protein dataset $\mathcal{D}$, radius $r_{\text{max}}=5$Å
$\mathcal{E}\leftarrow\emptyset$
@@ Appendix E Environment construction and AFM embedding
9:                       $\mathcal{X}_{a}\leftarrow\mathcal{X}_{a}\cup\{a^{\prime}\}$triangleright All atoms in $a^{\prime}$
@@
Appendix F Predicting amino acid, secondary structure, and pKa
@@
F.1 Secondary Structure and Amino Acid Prediction
@@
  • In amino acid prediction, the target variable $\mathrm{y}$ holds one of the following categories: $\{A,C,D,E,F,G,H,I,K,L,M,N,P,Q,R,S,T,V,W,Y\}$, each representing a standard twenty amino acids (cf. [32] for further explanation).
@@
[…]
@@
[…]
F.2 Acid dissociation constant (pKa) prediction
@@
where $N$ is the number of samples in the dataset, $y^{*}_{i}$ is the groundtruth pKa value of a protonated site in a protein obtained by solving the Poisson-Boltzmann equation from [55], and $\hat{y}_{i}$ is the predicted pKa.
Evaluation metrics.
F.3 Model Architecture
Across all three tasks, we used a consistent model architecture with little changes in hyperparameters. To incorporate the spatial and relation information between atomistic features, we constructed a fully connected graph between the atoms of the same residue. We propose to use a graph convolution network (GCN) [34] to predict the specific class type. The update rule at layer $l\in\mathbb{N}$ is defined as,
@@ F.3 Model Architecture
Here, $\bar{\mathbf{A}}=(\mathbf{I}+\mathbf{A})$ is the normalized adjacency matrix that serves as a low-pass filter when aggregating information over the neighbors. $\mathbf{A}$ is the graph’s adjacency matrix, and $\mathbf{I}$ is an identity matrix. Also, $\mathbf{W}^{(l)}$ is the learnable weight matrix to transform the features at layer $l$, $\mathbf{X}$ is the node feature matrix defined for layer $l$ of the GNN. We used SiLU (Sigmoid Linear Unit) [16] as the non-linear activation function and we applied layer normalization after each graph convolution operation to stabilize training. The operation in Eq. 12 is repeated for $L=10$ layers. The initial input $\mathbf{X}^{(0)}$ consists of the atomistic features for each atom. At the final layer ($l=L-1$), in the case of classification, the network outputs a vector of shape $|\mathrm{y}|$ representing class-wise logits, and the network outputs a scalar for regression tasks. Lastly, we use dropout [65] for every intermediate layer to prevent overfitting. Below is the table of hyper-parameters used for these experiments.
@@
F.4 Computation details

### v3

@@ B Isolated $\beta$-hridge
Table B.15: Set of protein structures with multi-modal backbone conformations (altlocs) from Rosenberg et al. (2024) guided using MLFF descriptors.
PDB ID Residue range Sequence Resolution [Å]
3V3S:B $\texttt{245}-\texttt{250}$ KAQERD $1.90$
7EC8:A $\texttt{187}-\texttt{190}$ DGGI $1.35$
5G51:A $\texttt{290}-\texttt{295}$ GSASDQ $1.45$
4NPU:B $\texttt{133}-\texttt{136}$ FEEI $1.50$
Appendix D Extended Background
@@
Chemical shifts. The resonance frequency of a nuclear spin, $\nu$, depends directly on the splitting of its energy levels in a magnetic field. This is given as $\nu\propto\gamma B$, where $B$ is the strength of the magnetic field that the nucleus experiences and $\gamma$ is the gyromagnetic ratio of the nucleus, a characteristic of an atom’s nucleus (e.g., 1H, 13C). In Nuclear Magnetic Resonance (NMR) experiments, $B$ is primarily determined by the externally applied magnetic field $B_{0}$, but at the level of an observed atom it is slightly perturbed by local magnetic fields $\Delta B$ induced by the surrounding electron cloud. These perturbations, which are typically several orders of magnitude weaker than $B_{0}$, cause small shifts in the resonance frequency – known as chemical shifts. The chemical shift reflects subtle variations in the local electronic environment and is therefore highly informative about molecular structure and composition. The chemical change of an atom is often reported in relative terms as $\delta=10^{6}\cdot\frac{\nu-\nu_{0}}{\nu_{0}}$, where $\nu$ is the observed resonance frequency and $\nu_{0}$ is the resonance frequency of the same type of nucleus in a reference compound – commonly the 1H signal of 2,2-dimethyl-2-silapentane-5-sulfonate (DSS) is used for the 1H chemical shift scale. Although dimensionless, chemical shifts $\delta$ are conventionally expressed in parts per million (ppm), hence the $10^{6}$ scaling factor in the definition.
@@
The electronic environment that gives rise to chemical shifts is influenced by factors such as the nature of chemical bonds, the identities of neighboring atoms, and their three-dimensional spatial arrangement. Additionally, dihedral angles [^3] The torsion angle about the bond between atoms B and C in a four-atom segment A-B-C-D play a significant role in modulating electron density around nuclei and thereby affect chemical shifts. Other factors that influence the chemical shift include hydrogen bonding interactions, ligand binding, or proximity to solvent molecules. Therefore, chemical shifts provide site-specific information about the local environment at the atomic-level. The chemical shift reports the time-averaged environment over all motions on time scales shorter than milliseconds. Chemical shifts are particularly valuable for inferring atomic connectivity, identifying functional groups, and detecting conformational changes, due to their sensitivity to fine variations in electronic structure (Claridge, 2016; Günther, 2013). Accurate chemical shift prediction plays a critical role in applications like automated NMR resonance assignment, molecular structure determination and validation, and the analysis of complex chemical systems (Wishart, 2011; Bermel et al., 2015). Despite this, reliable chemical shift prediction remains a challenge due to the sensitivity of chemical shifts to nuanced changes in molecular environment (Kuprov et al., 2007; Case, 2013).
Acid dissociation constants (pKa) pKa is a thermodynamic quantity describing the tendency of a titratable group to donate a proton. When the solvent pH is lower than the pKa, the group will be predominantly protonated; conversely, it becomes increasingly deprotonated at pH values above the pKa. Each amino acid has a canonical pKa in isolation (e.g., Aspartate $\sim$3.9), but within proteins these values can shift significantly. Nearby charged or polar residues, hydrogen-bond partners, and desolvation effects (especially in buried regions) modulate proton affinity and can shift pKa by several pH units. Accurate prediction of these shifts is crucial for understanding enzymatic catalysis, pH-dependent conformational changes, and protein stability.
Molecular Dynamics and Density Functional Theory.
Molecular Dynamics (MD) simulations numerically integrate Newton’s equations of motion to evolve the atomic positions of a system over time, requiring potential energies and forces at every timestep. When computed from quantum mechanics, these quantities can be obtained using Density Functional Theory (DFT), a first-principles electronic-structure method that solves the Kohn-Sham equations to deliver accurate total energies, forces, and electronic properties. However, the steep computational cost of DFT limits its direct application to small systems and short timescales. Machine learning force fields (MLFFs) address this limitation by providing near-DFT quality energies and forces at orders-of-magnitude lower cost, enabling quantum-accurate MD simulations of systems with tens of thousands of atoms over nanosecond timescales.
@@
ESM-based descriptors (Lin et al., 2023; Ouyang-Zhang et al., 2025; Hayes et al., 2025). Evolutionary Scale Modeling (ESM) is a family of transformer-based protein language models trained on a large-scale dataset of protein amino acid sequences using masked language modeling, similar to BERT (Devlin et al., 2019) for natural language. Unlike traditional methods that rely on multiple sequence alignments (MSAs) or co-evolutionary profiles, ESM models learn context-aware representations directly from amino acid sequences, enabling them to capture biochemical, structural, and evolutionary patterns within a protein sequence without explicit alignment. These representations have been used in two main ways. First, they underpin single-sequence 3D structure prediction in the ESMFold family of models (Lin et al., 2023). Second, later work augments these sequence-only models with structural supervision to obtain more structure-aware representations, for example in multimodal sequence-structure models such as ESM3 (Hayes et al., 2025) and in sequence-based protein language models that explicitly distill structural information into their embeddings (Ouyang-Zhang et al., 2025).
@@
MACE and Egret descriptors (Batatia et al., 2022; Wagen et al., 2025). Message passing neural networks (Gilmer et al., 2017; Schütt et al., 2017) typically exchange messages between pairs of nodes in a graph. In contrast, the MACE family of models (MACE, Egret) generalizes this framework by learning equivariant messages involving a system of $n$ (order parameter) nodes (or atoms) at a time. This enables richer geometric and chemical modeling by operating directly over higher-order interactions. Formally, for a given node at layer $t$ of the graph, MACE first learns the following $S_{n}$-invariant and $O(3)$-equivariant features between pairs of atoms:
@@
AIMNet2-based descriptors (Zubatyuk et al., 2019). AIMNet constructs Atomic Feature Vectors (AFVs) by iteratively embedding the local atomic environments and updating atomic representations using geometric information. Given the atomic coordinates $\mathbf{R}$ and atomic number $Z$, an initial atomic feature vector $\mathbf{A}_{i}^{(0)}$ is derived for atom $i$. The initial embeddings are parametrized as learnable vectors that depend on atomic number $Z_{i}$ of each atom.
@@
This process is repeated over multiple iterations, enabling information to propagate beyond the local neighborhood and capture non-local effects such as charge redistribution and polarization. After several refinement steps, AIMNet produces AFVs of fixed dimensionality ($256$), which are used to predict various molecular and atomic properties. Training is performed in a multitask setting, where the network jointly learns to predict a variety of molecular and atomic properties – including energies, forces, partial charges, and dipole moments – using shared representations to improve generalization. The model is trained on large datasets of quantum mechanical calculations, with regularization techniques to ensure stability and prevent overfitting.
OrbNet-based descriptors (Qiao et al., 2020). In similar spirit to semi-empirical molecular force-fields, OrbNet constructs feature vectors by leveraging symmetry-adapted atomic orbitals (AO), which capture local electronic structure information obtained from quantum calculations while respecting the symmetries inherent in molecular systems. Using mean-field quantum mechanics, OrbNet derives overlap matrices $\mathbf{S}$, Fock matrix $\mathbf{F}$, density matrix $\mathbf{D}$, and the Hamiltonian $\mathbf{H}$. Each matrix element $O^{\nu\mu}$ corresponds to interactions between atomic orbitals $\nu$ and $\mu$, and blocks $\mathbf{O}_{ij}$ represent interactions between atoms $i$ and $j$.
@@
LoCoHD (Fazekas et al., 2024) is a method for quantifying chemical and structural differences between protein environments. Unlike alignment-based or purely geometric measures, LoCoHD characterizes each local environment as a distribution of chemical “primitive types” such as atom types or residue centroids within a specified radius. The similarity between two environments is then measured using the Hellinger distance between their respective distributions. In this paper, to test the effectiveness of LoCoHD descriptors in representing local environments, we construct LocoHD embeddings using its representation of a local environment while computing the similarity metric. Each embedding is computed by aggregating statistics of the primitive types, weighted by their distance from the central residue. This results in a purely structural and chemical descriptor of the local environment.
Appendix E Data curation
The dataset used within the scope of this paper was sourced from RefDB (Zhang et al., 2003) which is a subset of the Biomolecular Magnetic Resonance Databank (BMRB) (Hoch et al., 2023). RefDB is a curated list of BMRB entries with calibrated chemical shifts. We select monomers from the RefDB and perform sequence redundancy filtering with mmseqs via,
mmseqs easy-cluster sequences.fasta --min-seq-id 0.5 -c 0.8 --cov-mode 5
which results in sequence clusters. We then split the clusters to define train and test sets. Our final dataset consists of $1048$ BMRB entries in RefDB ($225$ test + $823$ train), from where we collected the amino acid sequences, and the experimental chemical shifts.
Why use AlphaFold structures.
Instead of using Protein Data Bank (PDB) structures, we predict 3D structures for all entries with AlphaFold2 (Jumper et al., 2021) via OpenFold (Ahdritz et al., 2024). Using experimental crystal structures would introduce several limitations:
  • Poor generalization. X-ray crystallography covers only a small fraction of the proteome, meaning a model trained on PDB-derived environments would not generalize to most sequences without available structures.
  • Different experimental conditions. Crystal structures and NMR represent different physical states (solid vs. solution), leading to systematic differences in local geometry that can affect chemical shifts.
  • Sequence mismatches and ambiguity. We observed $\sim$15.8% sequence misalignment between BMRB entries and corresponding PDB structures. A single BMRB ID can map to multiple PDB entries (e.g., BMRB 15064 $\to$ 16 PDBs), making it unclear which structure should be used as reference.
AlphaFold alleviates these issues by providing systematic and consistent structural predictions for every sequence, including those with no experimental structure, enabling uniform training and evaluation.
Local protein environments are extracted from each AlphaFold structure as described in Algorithm 1.
Random-coil chemical shifts used in our predictor are obtained from UCBShift2 (Ptaszek et al., 2024).
Due to the limited availability of high-quality experimental training data, our pKa prediction models are regressed over synthetic values determined by solving the Poisson-Boltzmann equation via the method of Reis et al. (2020).
Appendix F Environment construction and MLFF embedding
This section elaborates on the amino acid centric environment construction algorithm discussed in Section 3.
Algorithm 1 Amino Acid Centric Environment Extraction
1:Protein dataset $\mathcal{D}$, radius $r_{\text{max}}=5$Å
@@ Appendix F Environment construction and MLFF embedding
9:        $\mathcal{X}_{a}\leftarrow\mathcal{X}_{a}\cup\{a^{\prime}\}$ $\triangleright$ All atoms in $a^{\prime}$
@@
F.1 Radius ablation
To select the radius used for environment construction, we performed an ablation study over $r\in\{4,5,6\}$ Å using MACE (Kovács et al., 2023) embeddings for backbone chemical shift prediction. MACE is trained with a cut-off of 6 Å, so larger radii are not meaningful. The results (Table B.9) show that a radius of 5 Å performs best. We hypothesize that this radius is large enough to capture the relevant NMR chemical-shift environment while still keeping most residue neighborhoods as complete as possible under the 6 Å cut-off, thereby providing a good balance between coverage and truncation.
F.2 MLFF Layer Comparison
This ablation study examines how selecting different hidden layers as MLFF embeddings affects performance for MACE and Orb. MACE has two hidden message-passing layers; throughout this work, the MACE descriptor is defined as the concatenation of the invariant components of their hidden states, matching the default behavior of the mace.get_descriptors function. For all other MLFF models, the descriptor corresponds to the hidden state of the final message passing layer. Orb has fifteen hidden message-passing layers, so in this case the descriptor is given by the hidden state of the 15th layer.
Table B.10 summarizes the per-atom mean absolute errors for all evaluated layers on the chemical-shift prediction task. For MACE, the concatenation of its two hidden layers yields the best performance for most atom types, while for Orb the 10th hidden layer performs best in most cases. Nevertheless, throughout this work we use the 15th Orb layer to maintain a consistent configuration across MLFFs with many layers, since concatenating all hidden states would result in impractically high-dimensional descriptors.
Appendix G Predicting amino acid, secondary structure, and pKa
@@
G.1 Secondary Structure and Amino Acid Prediction
@@
  • In amino acid prediction, the target variable $\mathrm{y}$ holds one of the following categories: $\{A,C,D,E,F,G,H,I,K,L,M,N,P,Q,R,S,T,V,W,Y\}$, each representing a standard twenty amino acids (cf. (Kawashima & Kanehisa, 2000) for further explanation).
@@
    $\text{F1}_{i}=2\cdot\dfrac{\text{Precision}_{i}\cdot\text{Recall}_{i}}{\text{Precision}_{i}+\text{Recall}_{i}}$
@@
We conduct an ablation study on the set of atoms $\mathcal{A}_{a}$ to evaluate the effect of including descriptors from multiple atoms within amino acid $a$ on secondary structure and amino acid prediction performance. Specifically, two configurations are compared: one where $\mathcal{A}_{a}=\{\text{C, CA, CB, N, H, HA}\}$, and a reduced version $\mathcal{A}_{a}=\{\text{CA}\}$. For this ablation, identical models are trained using Egret descriptors, varying only the atom set. Egret descriptors are used because as shown in Tables B.3 and B.4, they consistently outperform other atom-level descriptor types. The ablation results for secondary structure prediction are presented in Table B.13 and for amino acid prediction in Table B.14. Clearly, incorporating a larger set of atoms in $\mathcal{A}_{a}$ leads to a substantial improvement in prediction accuracy.
G.2 Acid dissociation constant (pKa) prediction
@@
where $N$ is the number of samples in the dataset, $y^{*}_{i}$ is the groundtruth pKa value of a protonated site in a protein obtained by solving the Poisson-Boltzmann equation from Reis et al. (2020), and $\hat{y}_{i}$ is the predicted pKa.
Comparison on experimental data.
To quantitatively assess the practical performance of our approach, we compare against pKa-ANI Gokcan & Isayev (2022) and PROPKA Olsson et al. (2011) on the benchmark summarized in Table B.7. Both reference methods were originally trained on experimental pKa values derived from PDB structures, in contrast our models were trained on simulated values and predominantly on AlphaFold2 structures.
For pKa-ANI, we evaluate on the PKAD-R (Ancona et al., 2023) dataset, a curated and updated version of the original PKAD dataset used for their training. Because no explicit train–test split was reported, we benchmark both models on all entries in PKAD-R. For PropKa, we evaluate on their training set, for which exact predictions were provided. The authors of PropKa also mention evaluation on a set of $201$ proteins assembled from three previously published datasets, but they do not specify the corresponding PDB identifiers or make their predictions available, preventing a direct comparison on that benchmark.
These comparisons are conservative with respect to our method, because the baselines are trained on experimental PDB-based data whereas our models are trained on simulated labels and AlphaFold structures. Nevertheless, our model outperforms PropKa on more than half of the entries and matches or surpasses pKa-ANI on $3$ out of $4$ residue types. For the remaining residue type (Histidine), we hypothesize that the weaker performance is due to the substantially smaller number of available training examples.
G.3 Model Architecture
Across all three tasks, we used a consistent model architecture with little changes in hyperparameters. To incorporate the spatial and relation information between atomistic features, we constructed a fully connected graph between the atoms of the same residue. We propose to use a graph convolution network (GCN) (Kipf & Welling, 2016) to predict the specific class type. The update rule at layer $l\in\mathbb{N}$ is defined as,
@@ G.3 Model Architecture
Here, $\bar{\mathbf{A}}=(\mathbf{I}+\mathbf{A})$ is the normalized adjacency matrix that serves as a low-pass filter when aggregating information over the neighbors. $\mathbf{A}$ is the graph’s adjacency matrix, and $\mathbf{I}$ is an identity matrix. Also, $\mathbf{W}^{(l)}$ is the learnable weight matrix to transform the features at layer $l$, $\mathbf{X}$ is the node feature matrix defined for layer $l$ of the GNN. We used SiLU (Sigmoid Linear Unit) (Elfwing et al., 2018) as the non-linear activation function and we applied layer normalization after each graph convolution operation to stabilize training. The operation in Eq. 12 is repeated for $L=10$ layers. The initial input $\mathbf{X}^{(0)}$ consists of the atomistic features for each atom. At the final layer (l=L1l=L-1), in the case of classification, the network outputs a vector of shape $|\mathrm{y}|$ representing class-wise logits, and the network outputs a scalar for regression tasks. Lastly, we use dropout (Srivastava et al., 2014) for every intermediate layer to prevent overfitting. Below is the table of hyper-parameters used for these experiments.
@@
Appendix H Using likelihoods to define environment similarity measure
H.1 Computation details